# Physics-informed Neural Networks for Heterogeneous Poroelastic Media

### v1

@@ 3.4.2 Hard enforcement of boundary/initial conditions
When the initial condition (IC) is “soft-enforced,” the deep neural network $\mathcal{N}(x,t,\theta)$ approximates the function $u(x,t)$ as $u(x,t)\approx\tilde{u}(x,t)=\mathcal{N}(x,t,\theta)$ with the initial condition of $u_{0}(x)$ at $t=0$. Conversely, in the case of “hard enforcement,” a new function $\tilde{u}$ is created to approximate $u(x,t)$ while satisfying the initial condition for all $\mathcal{N}(x,t,\theta)$, given by:
$\tilde{u}(x,t)=u_{0}(x)+t\mathcal{N}(x,t,\theta).$
For the “hard-enforcement” of the Dirichlet boundary condition, where $u_{D1}$ is imposed at $x=x_{\text{min}}$ and $u_{D2}$ at $x=x_{\text{max}}$, a function $\tilde{u}$ is created to approximate $u(x,t)$ as follows,
$\tilde{u}(x,t)=(x-x_{\text{min}})(x-x_{\text{max}})\mathcal{N}(x,t,\theta)-u_{D1}\frac{x-x_{\text{max}}}{x_{\text{max}}-x_{\text{min}}}+u_{D2}\frac{x-x_{\text{min}}}{x_{\text{max}}-x_{\text{min}}}.$
In both scenarios, there is no explicit need to specify the B/ICs in the loss functional, as the function $\tilde{u}(x,t)$ is meticulously crafted to automatically adhere to the relevant B/ICs.
@@ 4 Numerical Experiments
In this section, the performance of the two proposed frameworks, namely, I-PINNs with CoNN and I-PINNs with SNN, is compared with each other, as well as with conventional PINNs. This comparison is conducted on two benchmark examples examining poroelasticity in heterogeneous materials with a single interface. In Section  4.4, a comparative analysis is presented by solving the same benchmark problems with and without Glorot initialization and hard enforcement of B/ICs. To rigorously evaluate the accuracy of the proposed framework in addressing poroelastic problems, it’s imperative to compare the approximations with closed-form analytical solutions. Although such solutions are scarce for heterogeneous poroelastic scenarios, a few have been found for 1D problems [55]. The results of this study were directly compared with the exact solutions provided in this literature. All pertinent details, including the problem domain, interface location, material properties, and the closed-form analytical solutions for both problems, are extracted from Bean and Yi [55].
It is important to note that while the exact solutions conform to the boundary conditions (Eqs. (6a) and (6b)), they no longer satisfy the initial condition (Eq. (6c)). The initial condition is determined by evaluating the given exact solutions at $t=0$. For conciseness, the collectively proposed methods are referred to as I-PINNs (CoNN and SNN), while the conventional PINNs method is denoted simply as PINNs. The comparisons are based on two key factors: computational cost (total time for training the models) and accuracy, measured by the root mean square error against the closed-form analytical solution. Furthermore, both frameworks are enhanced with various modifications, as described in the previous section, and the impact of each modification on the accuracy of approximations is assessed. Finally, the convergence profiles of the CoNN and SNN models for the test cases are investigated.
Firstly, certain standards, such as the neural network architecture, total number of iterations, parameter scale for initialization, etc., must be established to ensure consistent comparisons. To determine these standards, we begin by considering a 1D homogeneous problem involving a compressible fluid. The material constants and the analytical solution for this problem are provided below -
@@
From Figure 6, it is evident that with a single layer and 5 hidden neurons, the RMSE is high for both displacement and pressure. As the number of neurons increases to 20, the root mean square rrror (RMSE) decreases, but further increments result in an upsurge. Analyzing the displacement plot reveals that this issue is promptly addressed by incorporating multiple hidden layers. With multiple hidden layers, even using only 5 hidden neurons in each layer results in low RMSEs. However, for pressures, the task is more challenging, necessitating a higher number of neurons in each layer and more hidden layers to reduce the RMSE. Given that the aim is to deal with coupled poroelasticity problems, it is very important to choose a specific layer architecture that consistently results in relatively low errors in approximating both output fields – displacements and pressures. In this regard, while a 4 hidden layer architecture with 10 neurons each may yield a lower RMSE for displacements (as seen in Figure 6(a)), it performs poorly for the pressure field (indicated in Figure 6(b)). Therefore, it is essential to choose a layer architecture that minimizes the RMSE for both output fields rather than optimizing for just one. Considering both output fields, only two combinations – 3 layers with 40 neurons each and 2 hidden layers with 80 neurons each – meet this criterion. We opt for the former among the two as the standard layer architecture for all subsequent experiments due to its relative simplicity, respecting computational cost considerations. The approximated pressure and displacement plots, along with their corresponding absolute error contour plots, are presented in Figure 7. For this simulation, we employed the aforementioned layer architecture consisting of 3 layers with 40 neurons each. It is evident that this specific layer architecture has effectively approximated both fields, with errors mostly lying close to $t=1$ (near the characteristic time).
@@ 4.1 Problem 1: Heterogeneous Poroelastic Media with an Incompressible Fluid
Training time 631 seconds 706 seconds 465 seconds
@@ 4.2 Problem 2: Heterogeneous Poroelastic Media with a Compressible Fluid
Training time 718 seconds 754 seconds 487 seconds
@@ 5 Conclusions
In this study, a novel Physics-Informed Neural Networks (PINNs) framework was introduced for modeling coupled problems specifically tailored for thermo-hydro-poromechanical systems in heterogeneous subsurface media. Notably, the framework’s application was limited to heterogeneous poroelastic media for this study, and the results showcased superior accuracy and convergence compared to the conventional PINNs framework. The approach incorporated two significant modifications to the neural network architecture. Firstly, a composite neural network (CoNN) was employed, where each sub-domain featured a dedicated neural network for each output field variable (displacements and pressures). These neural networks utilized the same activation functions but were trained separately for their network parameters (weights and biases). Secondly, the neural network was integrated into the Interface-PINNs (I-PINNs) framework, employing different activation functions across material interfaces to capture discontinuities in solution fields and gradients accurately. Collocation points for training the I-PINNs were prescribed using a biased sampling approach, prioritizing early-time solutions in accordance with causality. Additional modifications were introduced into the proposed framework, including initializing the network parameters using the Glorot scheme and hard enforcing the initial and boundary conditions. The accuracy and convergence behavior of the proposed framework was demonstrated through two specific numerical examples involving an incompressible fluid and a compressible fluid. A single neural network (SNN) architecture was also explored, integrated into the I-PINNs framework, and compared against the proposed CoNN. In comparison to the conventional PINNs framework, both proposed I-PINNs frameworks (CoNN and SNN) exhibited superior accuracy in benchmark numerical examples. While I-PINNs with SNN showed improved results over conventional PINNs, the standout performer was I-PINNs with CoNN, demonstrating better convergence behavior and improved accuracy, especially in the incompressible fluid example. Moreover, the training times for I-PINNs with SNN were slightly higher than those of I-PINNs with CoNN, making them computationally costlier. Moreover, it is essential to highlight the framework’s potential for extension to include thermo-mechanical aspects, which were not explicitly addressed in the current study. Looking ahead, future work involves expanding the proposed approach to address two-dimensional and three-dimensional poroelastic problems and further enhance its capabilities to model coupled multi-physics processes, including thermo-hydro-poromechanical problems involving composite materials or highly heterogeneous subsurface systems.

### v3

@@ 3.4.2 Hard enforcement of boundary/initial conditions
When the initial condition (IC) is “soft-enforced,” the deep neural network $\mathcal{N}(\mathbf{x},\theta)$ approximates the function $u(\mathbf{x})$ as $u(\mathbf{x})\approx\tilde{u}(\mathbf{x})=\mathcal{N}(\mathbf{x}\theta)$ with the initial condition of $u_{0}(\mathbf{x})$ at $t=0$. Conversely, in the case of “hard enforcement,” a new function $\tilde{u}$ is created to approximate $u(x,t)$ while satisfying the initial condition for all $\mathcal{N}(\mathbf{x},\theta)$, given by:
$\tilde{u}(\mathbf{x})=u_{0}(\mathbf{x})+t\mathcal{N}(\mathbf{x},\theta).$
For the “hard-enforcement” of the Dirichlet boundary condition, where $u_{D1}$ is imposed at $x=x_{\text{min}}$ and $u_{D2}$ at $x=x_{\text{max}}$, a function $\tilde{u}$ is created to approximate $u(\mathbf{x})$ as follows,
$\tilde{u}(\mathbf{x})=(x-x_{\text{min}})(x-x_{\text{max}})\mathcal{N}(\mathbf{x},\theta)-u_{D1}\frac{x-x_{\text{max}}}{x_{\text{max}}-x_{\text{min}}}+u_{D2}\frac{x-x_{\text{min}}}{x_{\text{max}}-x_{\text{min}}}.$
In both scenarios, there is no explicit need to specify the B/ICs in the loss functional, as the function $\tilde{u}(\mathbf{x})$ automatically satisfies the corresponding B/ICs.
@@ 4 Numerical Experiments
In this section, the performance of the two proposed frameworks, namely, I-PINNs with CoNN and I-PINNs with SNN, is compared with each other, as well as with conventional PINNs. This comparison is conducted on two benchmark examples examining poroelasticity in heterogeneous materials with a single interface. In Section  4.5, a comparative analysis is presented by solving the same benchmark problems with and without Glorot initialization and hard enforcement of B/ICs. To rigorously evaluate the accuracy of the proposed framework in addressing poroelastic problems, it’s imperative to compare the approximations with closed-form analytical solutions. Although such solutions are scarce for heterogeneous poroelastic scenarios, a few have been found for 1D problems [52]. The results of this study were directly compared with the exact solutions provided in this literature. All pertinent details, including the problem domain, interface location, material properties, and the closed-form analytical solutions for both problems, are extracted from Bean and Yi [52].
It is important to note that while the exact solutions conform to the boundary conditions (Eqs. (6a) and (6b)), they no longer satisfy the initial condition (Eq. (6c)). The initial condition is determined by evaluating the given exact solutions at $t=0$. For conciseness, the collectively proposed methods are referred to as I-PINNs (CoNN and SNN), while the conventional PINNs method is denoted simply as PINNs. The comparisons are based on two key factors: computational cost i.e. run-time for training the models and accuracy. The root mean square error (RMSE) metric is used to compare the numerical accuracy of the experiments, while for quantifying the relative computational cost of an experiment, we define the metric cost as: $C=t_{m}/t_{P}$, where $t_{m}$ represents the total training time for the method under consideration, and $t_{P}$ represents the total training time for the proposed method.
Furthermore, both frameworks are enhanced with various modifications, as described in the previous section, and the impact of each modification on the accuracy of approximations is assessed. Finally, the convergence profiles of the CoNN and SNN models for the test cases are investigated.
4.1 Hyperparameter-tuning:
Firstly, certain standards, such as the neural network architecture, total number of iterations, parameter scale for initialization, etc., must be established to ensure consistent comparisons. To determine these standards, the hyper-parameters are tuned on a simple 1D homogeneous problem involving a compressible fluid. The material constants and the analytical solution for this problem are provided below -
@@ 4.1 Hyperparameter-tuning:
From Figure 6, it is evident that with a single layer and 5 hidden neurons, the RMSE is high for both displacement and pressure. As the number of neurons increases to 20, the root mean square rrror (RMSE) decreases, but further increments result in an upsurge. Analyzing the displacement plot reveals that this issue is promptly addressed by incorporating multiple hidden layers. With multiple hidden layers, even using only 5 hidden neurons in each layer results in low RMSEs. However, for pressures, the task is more challenging, necessitating a higher number of neurons in each layer and more hidden layers to reduce the RMSE. Given that the aim is to deal with coupled poroelasticity problems, it is very important to choose a specific layer architecture that consistently results in relatively low errors in approximating both output fields – displacements and pressures. In this regard, while a 4 hidden layer architecture with 10 neurons each may yield a lower RMSE for displacements (as seen in Figure 6(a)), it performs poorly for the pressure field (indicated in Figure 6(b)). Therefore, it is essential to choose a layer architecture that minimizes the RMSE for both output fields, rather than optimizing for just one. Among the options, two architectures meet this criterion: 3 layers with 40 neurons each and 2 layers with 80 neurons each. The 3-layer configuration has a total of 6,804 parameters, whereas the 2-layer configuration has 13,602 parameters and is more computationally intensive. For practical purposes, the 3-layer architecture with 40 neurons each is selected for all subsequent experiments due to its simplicity and lower computational cost.
The approximated pressure and displacement plots, along with their corresponding absolute error contour plots, are presented in Figure 7. For this simulation, we employed the aforementioned layer architecture consisting of 3 layers with 40 neurons each. It is evident that this specific layer architecture has effectively approximated both fields, with errors mostly lying close to $t=1$ (near the characteristic time).
@@ 4.2 Problem 1: Heterogeneous Poroelastic Media with an Incompressible Fluid
Cost 1 1.2 0.8
@@ 4.3 Problem 2: Heterogeneous Poroelastic Media with a Compressible Fluid
Cost 1.0 1.2 0.7
@@ 4.5 Improvements on the Framework
5 Comparison with eXtended PINNs (XPINNs)
In this section, a series of 1D experiments are conducted (using the same problems as in the previous section) to compare the performance of the proposed method with a widely used domain-decomposition-based PINNs framework known as eXtended PINNs (XPINNs) [26] in terms of computational accuracy and cost. To ensure a fair comparison, the hyperparameters for both methods were selected to be as similar as possible. It is important to note that the key distinction between I-PINNs and XPINNs lies in their approach to subdomain decomposition along the interface: I-PINNs share the same set of parameters across subdomains, while XPINNs employ a different set of parameters for each neural network in the subdomains. Additionally, the loss terms related to the interface for the two methods are different. The proposed method enforces interface compatibility constraints as per Eqs. 13(h)-13(k), whereas the XPINNs implementation enforces flux continuity and the average value of the jump in the primary variable at the interface, as described in ‘Remark 5’ of Jagtap et al. [26]. Furthermore, to maintain a consistent bias–variance trade-off between the models, the number of training points in XPINNs was proportionately increased compared to I-PINNs
5.1 Comparison 1: Heterogeneous Poroelastic Media with an Incompressible Fluid
Firstly, the problem described in Section 4.2 is revisited, which involves incompressible fluid flow through a poroelastic medium with a material interface at $x=\zeta=1/5$, dividing the domain into two non-overlapping subdomains, $x_{1}\in(0,\zeta)$, and $x_{2}\in(\zeta,1)$. The two frameworks, I-PINNs–with–CoNN and XPINNs–with–CoNN, are applied using the hyperparameters listed in Table 4. From the outset, it is evident that although XPINNs utilize approximately 25% more parameters than I-PINNs, the root-mean-square error (RMSE) for both pressure and displacement approximations is of the order $\mathcal{O}(10^{-2})$, which is one order higher than that of I-PINNs ($\mathcal{O}(10^{-3})$ for both). Additionally, for this particular example, the computational cost of XPINNs is 40% higher than that of I-PINNs, indicating a direct relationship between computational cost and the number of trainable parameters.
Table 4: Hyper-parameters used in training I-PINNs and XPINNs models for approximating the solution to the incompressible fluid problem.
Parameter I-PINNs (CoNN) XPINNs (CoNN)
Layer architecture [2,20,20,20,1]$\times$2 [2,15,15,15,2]$\times$2
Activation functions GELU-tanh GELU-tanh
Number of training points [domain interior, external boundaries, interface] [4900,140,70] [5300,180,95]
Total # of parameters 1842 2164
Maximum iterations 20000 20000
RMSE (displacements) 3.8×$\times 10^{-3}$ 9.2×$\times 10^{-1}$
RMSE (pressures) 3.6×$\times 10^{-3}$ 1.5×$\times 10^{-2}$
Cost 1 1.5
5.2 Comparison 2: Heterogeneous Poroelastic Media with a Compressible Fluid
In the second experiment, the problem defined in Section 4.3 is considered, involving compressible fluid flow within a poroelastic medium with a material interface at $x=\zeta=1/3$. Both I-PINNs–with–CoNN and XPINNs–with–CoNN are utilized to approximate the solution, using the listed in Table 5. Similar to the previous example, XPINNs produced the greater RMSE of the two, with pressure errors being one order of magnitude higher and displacement errors being two orders of magnitude higher than those of I-PINNs. While these differences might appear minor in simple academic problems, they could become more significant in complex practical applications, especially those involving multiple interfaces or inverse problems.
Table 5: Hyper-parameters used in training I-PINNs and XPINNs models for approximating the solution to the compressible fluid problem.
Parameter I-PINNs (CoNN) XPINNs (CoNN)
Layer architecture [2,20,20,20,1]$\times$2 [2,15,15,15,2]$\times$2
Activation functions GELU-tanh GELU-tanh
Number of training points [domain interior, external boundaries, interface] [4900,140,70] [5300,180,95]
Total # of parameters 1842 2164
Maximum iterations 20000 20000
RMSE (displacements) 7.4×$\times 10^{-3}$ 1.4×$\times 10^{-2}$
RMSE (pressures) 8.1×$\times 10^{-3}$ 1.1×$\times 10^{-2}$
Cost 1 1.4
6 Discussions
While the approach presented in this manuscript has demonstrated promise in addressing coupled-field problems such as poroelasticity in heterogeneous media, several improvements to the approach are necessary before it can be applied to model field-scale applications. For instance, the numerical examples and comparative studies presented were restricted to 1D cases with material coefficient mismatches of $\mathcal{O}(10^{1})$. Looking ahead, the extension of the proposed method to two-dimensional and three-dimensional poroelastic problems, with complex topologies and diverse interface shapes, is of interest and will be the subject of future research. Furthermore, applications involving moving interfaces and greater mismatches in material parameters are also of significant interest and will be explored in subsequent studies. Moreover, the hyperparameters were selected using a manual grid search method; however, it is recommended that practitioners employ automatic hyperparameter optimization techniques, such as Bayesian optimization, to identify the optimal parameters and further enhance the framework’s performance. Although causality was enforced using a biased sampling approach, more advanced techniques like the causal training approach [53], or a time-window approach with exact enforcement of temporal continuity should be used for long-term prediction of time-dependent problems [54].
@@ 7 Conclusions
In this study, a novel Physics-Informed Neural Networks (PINNs) framework is introduced for modelling poroelasticity in heterogeneous media with material interfaces. The approach incorporates two modifications to the conventional PINNs architecture. Firstly, a composite neural network (CoNN) is employed, where each sub-domain features a dedicated neural network for each output field variable (displacements and pressures). These neural networks utilize the same activation functions but are trained separately for their network parameters (weights and biases). Secondly, the neural network is integrated into the Interface-PINNs or I-PINNs framework (Sarma et al. 2024; https://dx.doi.org/10.1016/j.cma.2024.117135), employing different activation functions across material interfaces to capture discontinuities in solution fields and gradients accurately. A single neural network (SNN) architecture integrated into the I-PINNs framework is also explored and compared against the proposed CoNN with respect to its accuracy and cost through two specific numerical examples involving an incompressible fluid and a compressible fluid.
For both examples, the root-mean-square-error (RMSE) of conventional PINNs is found to be at least two orders of magnitude worse than that of CoNN and SNN. Although CoNN and SNN yield comparable accuracies, the latter is $20\%$ more expensive than the former in their run-time. Moreover, a comparison is performed between I-PINNs (with CoNN) with another popular domain-decomposition method called eXtended PINNs (XPINNs (with CoNN)), for the same set of problems discussed above. The results demonstrate that XPINNs (with CoNN), despite being at least 40 percent more expensive in their runtimes, yield RMSEs at least one order of magnitude worse than I-PINNs (with CoNN).